### v1

# mO Revanth[#1]

Abstract
The growing collaboration between humans and AI models in generative tasks has introduced new challenges in distinguishing between human-written, AI-generated, and human-AI collaborative texts. In this work, we collect a multilingual, multi-domain, multi-generator dataset FAIDSet. We further introduce a fine-grained detection framework FAID to classify text into these three categories, meanwhile identifying the underlying AI model family. Unlike existing binary classifiers, FAID is built to capture both authorship and model-specific characteristics. Our method combines multi-level contrastive learning with multi-task auxiliary classification to learn subtle stylistic cues. By modeling AI families as distinct stylistic entities, FAID offers improved interpretability. We incorporate an adaptation to address distributional shifts without retraining for unseen data. Experimental results demonstrate that FAID outperforms several baseline approaches, particularly enhancing the generalization accuracy on unseen domains and new AI models. It provide a potential solution for improving transparency and accountability in AI-assisted writing.
1 Introduction
AI has progressively evolved from an assistant tool to a creator or initiator, from helping polish papers and generate code to initiate proposals and draft essays, while humans increasingly serve as optimizers and reviewers. In such deeply collaborative human-AI settings, measuring human contribution becomes challenging, yet clarifying authorship is critical for accountability and transparency, particularly in educational and academic contexts Wang et al. (2025). This work aims to identify human involvement by distinguishing the origin of a given text: fully AI-generated, fully human-written, or collaboratively produced, in multilingual context.
Numerous studies have explored multilingual AI-generated text detection, but most focus either on binary detection: human vs. AI Wang et al. (2024d, b), or on fine-grained detection limited to English Abassy et al. (2024); Koike et al. (2024); Wang et al. (2023); Zhang et al. (2024a). Moreover, both struggle with generalization to unseen domains, languages, and new AI generators Wang et al. (2024c); Li et al. (2024).
Our work aims to bridge this gap by (i) collecting a multilingual, multi-domain and multi-generator dataset for fine-grained detection, i.e., identifying a text into three categories: AI-generated, human-written, and human-AI collaborative; and (ii) introducing a detection framework FAID to ameliorate poor generalization performance. [^1] Our code and data are accessible at github.com/ngocminhta/FAID.
Our dataset focuses on academic field including domains of research paper abstracts, student theses, and scholarly reports, and involve recent LLM families, e.g., GPT, Gemini, DeepSeek and Llama OpenAI (2024); Team et al. (2024); DeepSeek-AI (2024); Dubey et al. (2024).
FAID treats each AI model as a distinct author, learning specific features in the hidden space to differentiate different authors, instead of classifying based on hand-crafted stylistic features. This was achieved by optimizing a language encoder with multi-level author relationship (e.g., stylistic similarity between texts from the same LLM family is greater than the similarity between a human and an AI) to capture author-specific distinguishable signals of an input text based on contrastive learning, along with the task of fine-tuning a classifier to recognize its origin. This multitask learning process forces the encoder to reorganize the hidden space, representations of texts from the same authors distributed closer than from different authors.
Practically, given that generations produced by models developed by the same company tend to have similar writing styles due to similar model architecture, training data and training strategy (more details in Appendix C), we consider each AI family as an author. This can also help the detector acquire prior knowledge of future LLMs from the same company. Experiments show that FAID consistently outperforms other baseline detectors in both in-domain and out-of-domain evaluations. Our contributions can be summarized as follows:
  • We collect a new multilingual, multi-domain, multi-generator dataset for fine-grained AI-generated text detection with 84k examples.
  • We propose a detection framework FAID to improve generalization performance in unseen domains and generators by capturing subtle stylistic features in hidden representation.
  • Experimental results show that FAID outperforms baseline detectors, particularly over unseen domains and generators. Meanwhile, the nature of FAID allows us to assess the stylistic proximity of a given text to others with ground truth authorship in the database, implicitly enhancing interpretability with cues explaining why a text is assigned to a specific class.
2 Backgrounds
Advanced LLMs have changed how we create and refine text. Instead of originating every word, humans now more often act as post-editors or reviewers, stepping in after AI drafts the initial content. This collaborative workflow spans fields from academic publishing and journalism to education and social media, making hybrid human-AI texts the new norm Cheng et al. (2024); Lee et al. (2022).
Such a shift calls question into binary authorship detection: human vs. AI. Texts can be fully human-written, fully AI-generated, or collaborative text (e.g., human-written, AI-polished; human-initiated, AI-continued; AI-generated, human-edited; and deeply mixed text) Artemova et al. (2025). This demands the fine-granularity authorship detection, and even tracing back to a specific model family, to measure human contributions Hutson (2025). Then, clear attribution of AI involvement helps uphold honesty in research and preserve scientific integrity. Responding to these challenges, we collect a new dataset, and propose a detection framework which generalizes well to new domains, languages, and models, showing high accuracy and reliability.
2.1 Fine-grained AI-generated Text Datasets
Many prior studies have explored fine-grained AI-generated text datasets across various forms of human-AI collaboration as shown in Table 3, for example, MixSet (Zhang et al., 2024b), and Beemo (Artemova et al., 2025). More datasets and detailed information regarding the label space, and the coverage of domains, languages and LLMs for each are summarized in Table 3.
Despite these contributions, they have been limited to English. There is a noticeable gap in the availability of large-scale, fine-grained multilingual AI-generated text detection datasets. To bridge this gap, we collect a multilingual fine-grained AI-generated text detection dataset. It encompasses 84k texts generated by the latest LLMs and includes diverse forms of human-AI collaborative generations. This dataset can facilitate the development of more robust and generalizable detection models that are capable of handling complex multilingual collaborative scenarios.
2.2 Generalization of AI-generated Text Detection
Recent studies such as M4GT Wang et al. (2024c) and LLM-DetectAIve Abassy et al. (2024) have highlighted a persistent challenge for both binary and fine-grained AI-generated text detection: poor generalization to unseen domains, languages, and generators. Many detection methods show a significant drop in performance on out-of-distribution data, underscoring the difficulty of building robust detectors to deal with real-world scenarios and evolving LLMs outputs.
Various approaches have been proposed to improve generalization. OUTFOX Koike et al. (2024) leverages adversarial in-context learning, dynamically generating challenging examples to enhance robustness, but still faces limitations in domain transferability and computational efficiency. LLM-DetectAIve Abassy et al. (2024) adopts fine-grained multi-way classification and incorporates domain-adversarial training to reduce overfitting, yet its generalization to truly unseen domains and generators remains limited. Also, its current version lacks multilingual support. SeqXGPT Wang et al. (2023) focused on sentence-level detection by combining log-probability features with convolutional and self-attention mechanisms, which helps capture subtle mixed-content signals and improves generalization across input styles. However, its reliance on specific model features and limited semantic representation constrains adaptability to new generators and domains.
A recent representative work, the DeTeCtive framework Guo et al. (2024), introduces multi-level contrastive learning to better capture writing-style diversity and enhance generalizability, especially for out-of-distribution scenarios, though it primarily focuses on binary classification and does not fully address human-AI collaborative texts.
2.3 Contrastive Learning for AI-Generated Text Detection
Contrastive learning has been extensively used to improve the sentence representation by pulling semantically similar sentences closer and pushing dissimilar ones apart, reorganizing the hidden space of sentence representation by semantics. For example, a representative work SimCSE regarded an input sentence under dropout noise as its semantically similar counterpart (i.e., positive pairs), and trained the encoder to close their distance. Additionally, they natural language inference (NLI) pairs, treating entailment pairs as positives and contradiction pairs as hard negatives (Gao et al., 2021). For negative pairs with contradictory meanings, the embeddings were trained to be distant in the latent space. Similarly, DeCLUTR Giorgi et al. (2021) constructed positive pairs by extracting different spans from the same texts, and negative pairs are sampled from different texts.
We adopt this philosophy to reorganize the latent space based on writing styles, pulling human-written texts to cluster together while remaining distant from AI-generated ones. Analogous to semantic textual similarity tasks, where sentence similarity ranges from 0 to 5 to reflect varying degrees of semantic overlap, we incorporate ordinal regression into our framework to model the degree of human involvement in collaborative texts, ranging from 0 (fully AI-generated) to 1 (fully human-written).
DeTeCtive also took advantage of contrastive learning and was used in binary AI-generated text detection Guo et al. (2024). Based on a multi-task framework, DeTeCtive was trained to learn writing-style diversity using a multi-level contrastive loss, and an auxiliary task of classifying the source of a given text (human vs. AI) to capture distinguishable signals. For inference, the pipeline encoded the input text as a hidden vector and applied dense retrieval to match the cluster, depending on the stylistic similarity with previously indexed training feature database. Additionally, instead of retraining the model when faced with new data, they encoded new data with the trained encoder to obtain embeddings, and then added them to the feature database to augment. This largely improved the generalizability in unseen domains and new generators.
However, this approach only considers two types of text: human-written and AI-generated, without further consideration of human-AI collaborative texts. Our work fills this gap with the goal of improving generalization performance for fine-grained AI-generated text detection.
3 FAIDSet
To fill the gap of fewer publicly-available multilingual fine-grained AI-generated text datasets, we collected a multilingual, multi-domain and multi-generator AI-generated text dataset FAIDSet. It contains texts generated by LLMs, written by humans and collaborated by both humans and LLMs, resulting in a total of $\sim$84k examples. See Appendix B for more details about FAIDSet.
FAIDSet covers two domains: student theses and paper abstracts, where identifying authorship is critical, across two languages including Vietnamese and English. We collected student theses from the database of Hanoi University of Science and Technology, and paper abstracts from arXiv and Vietnam Journals Online [^2] https://vjol.info.vn/ (VJOL).
Models and Label Space. We used the following multilingual LLMs families to produce both AI and human-AI collaborative texts: GPT-4/4o, Llama-3.x, Gemini 2.x, and Deepseek V3/R1. Regarding the human-AI collaborative text, we include AI-polished, AI-continued, and AI-paraphrased where the models are requested to polish or paraphrase inputs while ensuring the accuracy of any figures and statistics.
Diverse Prompt Strategies. We generate data with diverse tones and contexts while ensuring the accuracy of the content. Depending on the data source and context, we craft prompts to create varied outputs suitable for different real-world application scenarios. We generate responses with different tones using prompts such as "You are an IT student…" and "…who are very familiar with abstract writing…", which are described more in Appendix B.2.
Quality Control. Given that the task of AI-generated text detection naturally has gold labels when collecting data, we perform quality control by randomly sampling 10-20 instances for each domain, source, and generative model in data collection, and manually assess whether or not the generated texts are fluent and logically reasonable, to avoid incoherence, disorganization, or illogical texts. Such cases introduce remarkably distinguishable signals from human-written text, affecting the data quality. We adjust either the prompts or the generation parameters to address these problems.
4 Methodology
In this section, we first clarify task definition and then elaborate our detection method.
4.1 Task Definition
In this work, we formulate the task as a three-class classification problem: human-written, AI-generated, and human-AI collaborative texts. The human-AI collaborative category involves a range of interaction between humans and AI systems, such as human-written, AI-polished; human-initiated, AI-continued; human-written, AI-paraphrased; AI-generated, AI-humanized, etc.
Given the growing variety and complexity (e.g., deeply mixed text) of collaborative patterns, we do not exhaust all patterns. Instead, we consolidate all forms of human-AI collaboration into a single label to maintain practical simplicity and model generalizability. This approach reflects real-world usage, where using AI tools to enhance clarity or expression is increasingly common and often ethically acceptable.
We also investigate the dataset and find that the AI models within the same family tend to have similar writing styles and text distributions, due to their shared training data and architecture (see Appendix C). Therefore, we consider each model’s family to be an author with unique writing styles.
4.2 Framework Overview
Figure 1: Training architecture. Leveraging multi-level contrastive learning loss, we fine-tune a language model (we select XLM-RoBERTa Vamvas and Sennrich (2023), see Appendix D) based on the human, human-AI and AI texts, to force the model to reorganize the hidden space, pulling the embeddings within the same author families closer, and pushing the embeddings from different authors farther. We train an encoder that can represent text with distinguishable signals to discern authorship of text.
We aim to enable the encoder to capture multi-level similarities between authors (AI, human and human-AI) as Eq (6), forcing model to learn features that can differentiate distinct authors in a high-dimensional hidden vector space. We denote $\mathrm{S}_{c}$ as cosine similarity, $\phi(\cdot)$ is the encoder function and $P_{i}$, $P_{j}$, ($51\leq i\leq j<$ 5) are distributions of different text subsets. We have:
$\begin{split}\mathbb{E}_{x\in P_{i},y\in P_{j}}&\left[\mathrm{S}_{c}\left(\phi(x),\phi(y)\right)\right]\\ &\geq\mathbb{E}_{x\in P_{i},y\in P_{j+1}}\left[\mathrm{S}_{c}\left(\phi(x),\phi(y)\right)\right]\end{split}$ (1)
where $P_{1}$ corresponds to the distribution generated by a particular LLM family, $P_{2}$ to the distribution generated by any LLMs, $P_{3}$ to the distribution of a collaborative text generated by humans and the LLM family of $P_{1}$, $P_{4}$ to the distribution of collaborative text generated by humans and any LLM families, and $P_{5}$ to the distribution of human-written text.
To clarify the rationale for configuring FAID to expect that the similarity of a text $x$ (from lower-level distributions $P_{1}$ or $P_{2}$) with samples from $P_{3}$ is generally greater than or equal to its similarity with samples from $P_{4}$, consider the following.
  • If $x\in P_{1}$ ($x$ is generated by a particular LLM): Let $y_{\textit{LHS}}$ be drawn from $P_{3}$ and $y_{\textit{RHS}}$ be drawn from $P_{4}$. Naturally, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ should be greater than $\mathrm{S}_{c}(x,y_{\textit{RHS}})$. This is because $P_{3}$ contains texts that share a direct LLM family origin with $x$.
  • If $x\in P_{2}$ ($x$ is generated by any LLM): In this scenario, with $y_{\textit{LHS}}$ from $P_{3}$ and $y_{\textit{RHS}}$ from $P_{4}$ as defined above, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ is generally expected to be equal to $\mathrm{S}_{c}(x,y_{\textit{RHS}})$. Since $x$ can originate from any LLM, it does not inherently possess a stronger connection to the specific LLM family in $P_{3}$ than to the broader human-LLM collaborations represented in $P_{4}$.
This configuration aims to ensure that closeness in distribution corresponds to heightened similarity after encoding, encouraging the model to discern fine-grained multi-level relations.
4.3 Multi-level Contrastive Learning
Given a dataset with $N$ samples, the $i^{th}$ sample is denoted as $T_{i}$ and assigned with three-level labels indicating its source.
  • $x$: if $T_{i}$ is fully AI-generated, $x_{i}=0$, otherwise $x_{i}=1$;
  • $y$: if $T_{i}$ is fully human-written, $y_{i}=0$, otherwise $y_{i}=1$;
  • $z$: an indicator of a specific LLM family.
The encoder $\phi(\cdot)$ represents the sample $T_{i}$ in a $d$-dimensional vector space $\mathbb{R}^{d}$. Then we calculate the cosine similarity between two samples $T_{i}$ and $T_{j}$, denoted by: $\sigma(i,j)=\mathrm{S}_{c}\left(\phi(T_{i}),\phi(T_{j})\right)$.
For AI-generated text ($x=0$), the similarity between $T_{i}$ and another AI-generated text $T_{j}$ is greater than that with a human-written or collaborative text $T_{k}$ ($x=1$):
$\sigma(i,j)>\sigma(i,k),\forall x_{i}=0,x_{i}=x_{j},x_{k}=1$ (2)
If $x=0$, then the text is fully AI-generated. For this case, we do not consider $y$ since AI-generated text is considered as non-AI-collaboration. We can imply that the similarity between two texts written by the same AI is higher than that of two different AI families. Hence:
$\sigma(i,j)>\sigma(i,k),\forall x_{i}=0,z_{i}=z_{j},z_{i}\neq z_{k}$ (3)
The reverse condition (with $y=1$) is also true. We can conclude that:
$\sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i}=y_{j},y_{i}\neq y_{k}$ (4)
For cases where all samples are human-AI collaborative texts ($x,y=1$), two texts with involvements of the same AI family tend to be more similar than the ones that involve contributions from different families. That is:
$\sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i,j,k}=1,z_{i}=z_{j}\neq z_{k}$ (5)
Combining all, the text representation is learned with the following constraints:
$\begin{cases}\sigma(i,j)>\sigma(i,k),\forall x_{i}=0,x_{i}=x_{j}\neq x_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=0,z_{i}=z_{j}\neq z_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=1,x_{i}=x_{j}\neq x_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i}=y_{j}\neq y_{k};\\ \begin{split}\sigma(i,j)>\sigma(i,k),\forall&x_{i}=1,y_{i,j,k}=1,\\ &z_{i}=z_{j},z_{i}\neq z_{k}\\ \end{split}\end{cases}$
To enforce the similarity constraints outlined in Eq (6), we build upon the SimCLR framework Chen et al. (2020) and introduce a strategy for defining both positive and negative sample pairs, which forms the basis of our contrastive learning loss. Departing from traditional contrastive losses that rely on a single positive sample, our approach considers a group of positive instances that satisfy specific criteria. The similarity between the anchor and positive samples is computed as the average similarity across this entire positive set. For negative samples, we follow the methodology used in SimCLR. The resulting contrastive loss, expressed in Eq (7), involves $q$ as the anchor sample, $K^{+}$ as the positive sample set, $K^{-}$ as the negative sample set, $\tau$ as the temperature parameter, and $N_{K^{+}}$ as the number of positive samples.
$\mathcal{L}_{q}=-\log\frac{\exp\left(\sum_{k\in K^{+}}\frac{S(q,k)}{\tau}/N_{K^{+}}\right)}{\exp\left(\sum_{k\in K^{+}}\frac{S(q,k)}{\tau}/N_{K^{+}}\right)+\sum_{k\in K^{-}}\exp\left(\frac{S(q,k)}{\tau}\right)}$
Different constraints result in varied sets of positive and negative samples. Based on these sets, contrastive losses are computed at multiple levels. As we declared in Eq (7), each inequality in Eq (6) is denoted as $\mathcal{L}_{q_{i},\varepsilon}$ where $\varepsilon=\overline{1,5}$ respectively. To form Eq (8), we need to add the coefficients $\alpha,\beta,\gamma,\delta$, and $\zeta$ to maintain the balance of multi-level relations.
$\begin{split}\mathcal{L}_{mcl}=&\sum\limits_{i=1}^{N}\left[x_{i}\left(\alpha\mathcal{L}_{q_{i},1}+\beta\mathcal{L}_{q_{i},2}\right)+\right.\\ &\left.\left(1-x_{i}\right)\left(\gamma\mathcal{L}_{q_{i},3}+\delta\mathcal{L}_{q_{i},4}+\zeta\mathcal{L}_{q_{i},5}\right)\right]\end{split}$ (8)
Therefore, due to the last inequality in Eq (6) only specifying a case ($y=1$), and the other cases considering both values for $y$, so we have $\zeta=2\gamma=2\delta$. Also, to maintain the equilibrium, we need to keep $\alpha+\beta=\gamma+\delta+\zeta$. We set $\gamma=\delta=1$, then $\zeta=2,\alpha=\beta=2$.
This approach encourages the model to capture subtle and detailed features from different sources. As a result, the model becomes more adept at recognizing variations in writing styles. This capability enhances its accuracy and strengthens its ability to generalize when detecting AI-generated text.
4.4 Multi-task Auxiliary Learning
Multi-task learning Caruana (1997) allows a model to learn several tasks concurrently by sharing relevant information across them. This joint learning process helps the model develop more general and distinctive features. As a result, it improves the model’s ability to generalize to new data. Building on the previously described contrastive learning framework, we extend the encoder by attaching an MLP classifier at its output layer. This classifier is responsible for binary classification, determining whether a given query text was written by a human or generated by a LLM. Denote that the probability of $i^{th}$ sample with label $x_{i}=0$ is $p_{i}$. To train this component, we apply a cross-entropy loss function, denoted as $\mathcal{L}_{ce}$, defined as follows:
$\mathcal{L}_{ce}=-\dfrac{1}{N}\sum\limits_{i=1}^{N}x_{i}\log p_{i}+\left(1-x_{i}\right)\log\left(1-p_{i}\right)$ (9)
Therefore, the overall loss is computed as:
$\mathcal{L}=\mathcal{L}_{ce}+\mathcal{L}_{mcl}$ (10)
Figure 2: Inference architecture. (a) embed the input text into embedding vector using the fine-tuned encoder; (b) use Fuzzy KNN algorithm to cluster, retrieving which cluster the input text belongs to (see more in Appendix E). (c) The stored vector database $\mathcal{VD}$ was created by saving all embeddings of texts in training and validation sets using the fine-tuned encoder. If the input text is unseen, we embed the unseen text and save into a temporary vector database $\mathcal{VD}^{\prime}$, enhancing the generalization of the detector.
4.5 Handling Unseen Data without Retraining
Unseen data, whether from an unseen domain or generated by an unfamiliar model, remains a significant challenge even for state-of-the-art AI-generated text detection methods, due to the increasing capabilities of large language models. We tried to use the model classifier on its own, but we ended up by using a vector database along with Fuzzy k-Nearest Neighbors, which is illustrated in Figure 2. The results were written in Appendix E. Specifically, when dealing with unseen data, we use our trained model to embed these samples and add them to our existing vector database. Through careful parameter tuning, this approach enables our system to effectively handle newly encountered unseen data without the need retraining.
5 Experiments
In this section, we first describe datasets and baselines used, followed by two experiments evaluating FAID in classifying a text between human, AI and human-AI; and in identifying specific generators.
Scenario Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
In-domain, Known generator FAIDSet LLM-DetectAIve 94.34 94.45 93.79 94.10 0.1888 0.1107
SVM (N-grams) 84.41 84.53 82.93 83.62 0.5630 0.2916
T5-Sentinel 93.31 94.92 93.10 93.15 0.2104 0.1101
FAID 95.58 95.78 95.33 95.54 0.1719 0.0875
LLM-DetectAIve LLM-DetectAIve 95.71 95.78 95.72 95.71 0.1606 0.1314
SVM (N-grams) 83.43 76.09 71.84 73.60 0.2878 0.2064
T5-Sentinel 94.77 94.70 92.60 93.60 0.1663 0.1503
FAID 96.99 95.29 88.14 91.58 0.1561 0.0754
HART LLM-DetectAIve 94.39 94.25 94.33 94.29 0.3244 0.1789
SVM (N-grams) 58.93 59.75 61.07 60.30 1.2334 0.6849
T5-Sentinel 86.68 87.25 87.69 87.38 0.4339 0.2334
FAID 96.73 97.61 98.05 97.80 0.4631 0.1806
Unseen domain, Unseen generator Unseen domain LLM-DetectAIve 52.83 47.31 64.62 53.28 0.4733 0.4722
SVM (N-grams) 39.28 43.13 29.41 22.77 0.7689 0.6611
T5-Sentinel 55.56 49.54 66.67 55.34 0.4444 0.4444
FAID 62.78 70.73 71.77 69.46 0.4514 0.4486
Unseen generators LLM-DetectAIve 75.71 73.25 75.63 74.30 0.3714 0.2957
SVM (N-grams) 74.19 61.88 50.17 55.17 0.4324 0.3162
T5-Sentinel 85.95 85.77 84.59 85.16 0.3648 0.2419
FAID 93.31 92.40 94.44 93.25 0.1691 0.1167
Unseen domain and Unseen generators LLM-DetectAIve 62.93 66.74 71.17 61.97 0.4479 0.3964
SVM (N-grams) 34.50 39.38 26.67 21.70 0.9186 0.7429
T5-Sentinel 57.07 49.82 66.61 55.45 0.4314 0.4300
FAID 66.55 74.44 73.57 72.58 0.3939 0.3167
Table 1: Performance with three labels. For in-domain and known generator datasets, we train detectors on training set and evaluate using test set. For unseen data evaluation, we use the detector trained on FAIDSet, and evaluate on the unseen datasets. The best results are in bold and the second best are underlined.
5.1 Datasets
In addition to FAIDSet, for in-domain evaluation, we used another two datasets:
LLM-DetectAIve Abassy et al. (2024) encompasses various domains including arXiv, Wikihow, Wikipedia, Reddit, student essays, and peer reviews. The original human-written and machine-generated texts were augmented using a variety of LLMs (e.g., Llama 3, Mixtral, Gemma) to create 236,000-example dataset with two new labels: machine-written then machine-humanized, and human-written then machine-polished.
HART Bao et al. (2025) is a dataset spans domains of student essays, arXiv abstracts, story writing, and news articles. It has four categories: human-written texts, LLMs-refined texts, AI-generated texts, and humanized AI-generated texts. They further generated 21,500 cases based on this dataset for unbalanced labels.
To evaluate generalization ability of FAID on unseen scenarios, we collected the following data:
Unseen domain: We create a dataset consisting of 150 IELTS Writing essays from Kaggle ibrahimmazlum (2023), where all original texts are human-written. These essays were then used to generate human-AI collaborative and AI-generated texts, using the same models with FAIDSet.
Unseen generators: We selected 150 human-written abstracts from the FAIDSet test set and generated data for remaining labels using the three new LLM families: Qwen 2, Mistral, and Gemma 2.
Unseen domain & generators: Based on the human-written IELTS essays above, instead of the original four LLMs, we used the same LLM families as Unseen generator test set to generate data for the AI and human-AI labels, following the same process as with the unseen domains.
Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$
FAIDSet LLM-DetectAIve 0.7596 0.7697 0.7690 0.7653
SVM (N-grams) 0.6764 0.6515 0.6126 0.6312
T5-Sentinel 0.7568 0.7985 0.7840 0.7837
FAID 0.7964 0.8328 0.8352 0.8327
LLM-DetectAIve LLM-DetectAIve 0.9049 0.9064 0.8352 0.8693
SVM (N-grams) 0.8567 0.6516 0.6127 0.6313
T5-Sentinel 0.8154 0.8137 0.8009 0.8105
FAID 0.9089 0.8817 0.8672 0.8737
HART LLM-DetectAIve 0.8900 0.8787 0.8674 0.8715
SVM (N-grams) 0.6347 0.5454 0.4385 0.4806
T5-Sentinel 0.7852 0.7713 0.7834 0.7759
FAID 0.8996 0.9157 0.8648 0.8667
Table 2: Accuracy of identifying generators: human, GPT, Gemini, Deepseek, and Llama. The best is in bold.
5.2 Baselines
LLM-DetectAIve:
Abassy et al. (2024) fine-tuned a RoBERTa Liu et al. (2019a) on fine-grained English AI-generated texts with learning rate of $2\times 10^{-5}$ for 10 epochs.
Support Vector Machine with N-grams:
Prova (2024) utilized SVM with N-gram features for binary AI text detection (Human vs. AI). We adapted this approach to a three-class detection. We first generated word-level trigram features using Count Vectorizer, and then used a Linear SVM to train with the gold labels to guide the classification.
T5-Sentinel:
Chen et al. (2023) utilized T5-Sentinel to classify text into two labels (Human vs. AI). This approach leverages the T5 model to predict the conditional probability of the next token, integrating the classification task into a sequence-to-sequence completion framework. This unique method directly leverages the strength of encoder-decoder model, which are trained to understand context and generate coherent text.
5.3 Human, AI or Human-AI?
Table 1 shows the performance of FAID and three baselines in four evaluation settings. FAID consistently achieves the best accuracy in (i) in-domain and known generators, (ii) unseen domain, (iii) unseen generators, and (iv) unseen domain & generators settings, followed by LLM-DetectAIve in (i) and (iv), and T5-Sentinel in (ii) and (iii). Taking advantage of pretrained language models (i.e., RoBERTa, T5), they outperform SVM in most cases, particularly on HART dataset.
FAID’s approach enhances the generalization performance over unseen domains and generators, compared with other methods. From the absolute accuracy, we can find that generalizing to (ii) unseen domains and (iv) unseen domain & generator remains challenging, with accuracy of 62.78% and 66.55% respectively. The result suggests that FAID is an effective method to address the multilingual fine-grained AI-generated text detection task, improving the performance by leveraging multi-level contrastive learning to capture generalizable stylistic differences tied to AI families, rather than overfitting to surface-level artifacts.
5.4 Identifying Different Generators
We further examine the detector to identify specific AI generators which were seen during training. The goal of FAID is not only to detect AI involvement but also to identify the specific AI model or family applied, treating them as distinct authors. As shown in Table 2, across the three datasets, FAID consistently achieves higher performance compared to other baselines in almost all metrics. LLM-DetectAIve achieves comparable results with our detector FAID on the test set of LLM-DetectAIve dataset, except that its precision is slightly lower. FAID’s high performance when dealing with text originating from diverse known generators within these datasets indicates that FAID learned unique writing patterns and features of different AI generators by leveraging multi-level contrastive learning.
6 Conclusion and Future Works
In this work, we collected a multilingual fine-grained AI-generated text detection dataset FAIDSet with $\sim$84,000 examples. We further introduced FAID, a fine-grained AI-generated text detection framework aiming to distinguish between human-written, AI-generated, and human-AI collaborative texts. It combines multi-level contrastive learning and multi-task auxiliary objectives, with the perspective of treating AI model families as stylistic authors. FAID captures subtle writing patterns that generalize effectively across domains and evolving AI systems, where Fuzzy KNN-based inference method and training-free incremental adaptation enhance the robustness, enabling strong performance on unseen data.
Experiments demonstrate that FAID consistently outperforms other strong baselines across multiple datasets and configurations. Its capability to recognize nuanced forms of collaborative texts and the ability to adapt to new generative models make it particularly suited for real-world applications, improving transparency in AI-assisted writing.
In the future, we plan to expand FAIDSet by incorporating more languages, generators and domains, especially low-resource languages and informal text genres such as social media content or student essays, to further enhance cross-lingual and domain robustness. We also continue to investigate to deal with AI-generated text or more complex human-AI collaborative texts (i.e., from two or more AI families, generated by multiple AI families then edited by human) to deal with human’s behaviors in using AI for writing.
7 Limitations
While FAID demonstrates strong performance and generalization across various domains and AI models, several limitations remain. First, although our dataset is multilingual and multi-domain, it is still limited in terms of low-resource languages and niche writing domains, which may affect performance in those contexts. Second, our framework is based on the fact that texts produced by AI models from the same family share similar stylistic features. However, this assumption may break down when a single text is influenced by multiple AI systems – such as when a human uses different models for drafting, rewriting, and polishing. In such cases, the resulting style may blend traits from multiple sources, making it more difficult to attribute authorship to a single AI family or clearly distinguish collaboration boundaries.
8 Ethics and Broader Impacts
Data Collection and Licenses
A primary ethical consideration is the data license. We reused existing dataset for our research: LLM-DetectAIve, HART and IELTS Writing, which have been publicly released under clear licensing agreements. We adhere to the intended usage of these dataset licenses.
Security Implications.
FAIDSet streamlines both the creation and rigorous testing of FAID. By spotting AI-generated material, FAID helps preserve academic integrity, flag potential misconduct, and protect the genuine contributions of authors. More broadly, it supports efforts to prevent the misuse of generative technologies in areas such as credential falsification. Detecting AI-crafted content across different languages can be tricky, due to language’s unique grammar and style. By enabling robust, multilingual and multi-generator detection with accurate results, FAIDSet empowers people everywhere, especially in academic scenarios, to deploy AI responsibly. At the same time, it fosters critical digital literacy, giving everyone a clear understanding of both the strengths and limits of generative AI.
Responsible Use of AI-generated Text Detection.
FAID is designed to enhance transparency in AI-assisted writing by enabling the fine-grained detection of AI involvement in text generation. While this has clear benefits for academic integrity, and content provenance, we acknowledge the potential for misuse. For instance, such tools could be used to unfairly penalize individuals in educational or professional settings based on incorrect or biased predictions. To mitigate this, we stress that FAID is not intended for high-stakes decision-making without human oversight.
Bias and Fairness.
AI-generated text detection systems may inadvertently encode or amplify biases present in training data. FAIDSet has been carefully constructed to include diverse domains and languages to reduce such biases. Nonetheless, we encourage ongoing auditing and benchmarking of fairness across populations and writing styles, and welcome community feedback for further improvements.
Transparency and Reproducibility.
In this work, to promote open research and community contributions, we published our code and data to the public.
Appendix
Appendix A Some current datasets on AI-generated texts
We do analysis in many current dataset on AI-generated text and summarized them in Table 3. Many dataset contains fine-grained labels and multi-generator, but all of analyzed datasets are monolingual – only in English. This problem makes us to collect and formulate a new dataset for multilingual, multi-domain and multilingual AI-generated text dataset to adapt the approach to real-world scenarios.
Dataset Languages Label Space Domains Generators Size
MixSet
Zhang et al. (2024c)
English
Human-written, AI-polished
Human-initiated, AI-continued
AI-written, human-edited
Deeply-mixed text
Email
News
Game reviews,
Paper abstracts,
Speech,
Blog
GPT-4
Llama 2
Dolly
3,600
DetectiAIve
Abassy et al. (2024)
English
Human-written
AI-generated
Human-written, AI-polished
AI-written, AI-humanized
arXiv abstracts,
Reddit posts,
Wikihow,
Wikipedia articles,
OUTFOX essays,
Peer reviews
GPT-4o
Mistral 7B
Llama 3.1 8B
Llama 3.1 70B
Gemini
Cohere
487,996
Beemo
Artemova et al. (2025)
English
AI-generated, AI-humanized
Human-written
AI-generated
AI-written, human-edited
Generation,
Rewrite,
Open QA,
Summarize,
Closed QA
Llama 2
Llama 3.1
GPT-4o
Zephyr
Mixtral
Tulu
Gemma
Mistral
19,256
M4GT
Wang et al. (2024c)
English
Human-written, AI-continued
Human-written
AI-generated
Peer review,
OUTFOX
Llama 2
GPT-4
GPT-3.5
33,912
Real or Fake
Dugan et al. (2022)
English
Human-written
Human-initiated, AI-continued
Recipes,
Presidential Speeches,
Short Stories,
New York Times
GPT-2,
GPT-2 XL
CTRL
9,148
RoFT-chatgpt
Kushnareva et al. (2024)
English Human-initiated, AI-continued
Short Stories,
Recipes,
New York Times,
Presidential Speeches
GPT-3.5-turbo 6,940
Co-author
Zeng et al. (2024)
English Deeply-mixed text
Creative writing,
New York Times
GPT-3 1,447
TriBERT
Zeng et al. (2023)
English
Human-initiated, AI-continued
Deeply-mixed text
Human-written
Essays ChatGPT 34,272
LAMP
Chakrabarty et al. (2025)
English AI-generated, human-edited Creative writing
GPT-4o
Claude 3.5 Sonnet
Llama 3.1 70B
1,282
APT-Eval
Saha and Feizi (2025)
English Human-written, AI-polished Based on MixSet
GPT-4o
Llama 3.1 70B
Llama 3.1 8B
Llama 2 7B
11,700
HART
Bao et al. (2025)
English
Human-written
Human-written, AI-polished
AI-generated, AI-humanized
AI-generated text
AI-generated, human-edted
GPT-3.5-turbo
GPT-4o
Claude 3.5 Sonnet
Gemini 1.5 Pro
Llama 3.3 70B
Qwen 2.5 72B
16,000
LLMDetect
Cheng et al. (2024)
English
Human-written
Human-written, AI-polished
Human-written, AI-extended
AI-generated
DeepSeek v2
Llama 3 70B
Claude 3.5 Sonnet
GPT-4o
64,304
ICNALE corpus
Macko et al. (2025)
English Human-written Essays
Qwen 2.5
Llama 3.1 8B/70B
Llama 3.2 1B/3B
Mistral Small
67,000
CyberHumanAI
Najjar et al. (2025)
English Human-written Wikipedia 500
Table 3: English Fine-grained AI-generated Text Detection Datasets Overview.
Appendix B FAIDSet Statistics and Analysis
B.1 Statistics
Our FAIDSet includes 83,350 examples, which are separated into three subsets: train, validation, and test with the ratio shown in Table 4. The data set also comprises various sources of human-written texts, which are described in Table 5.
Subset Human AI Human-AI collab.
Train 14,176 12,076 32,091
Validation 3,038 2,588 6,876
Test 6,879 2,599 3,038
Table 4: Number of examples per label in subsets in FAIDSet dataset.
Source No. Human texts
arXiv abstracts 2,000
VJOL abstracts 2,195
HUST theses (English) 4,898
HUST theses (Vietnamese) 11,164
Table 5: Statistics of human-written text’s origins in FAIDSet.
B.2 Diverse Prompt Strategies
To avoid biasing our generated corpora toward any single style or topic, we employ a broad set of prompt templates when synthesizing AI-generated and human-AI collaborative texts. By varying prompt structure, content domain, and complexity, we ensure that the resulting outputs cover a wide distribution of writing patterns, vocabulary, and rhetorical devices. This diversity helps our detector generalize more effectively to real-world inputs.
Concretely, we sample two out of five prompts for AI-generated texts and several categories of human-AI collaborative texts in Table 6:
  • AI-polished: Texts that were originally written by a human and then lightly refined by an AI system to improve grammar, clarity, or fluency without altering the core content or intent.
  • AI-continued: Texts where a human wrote an initial portion (e.g., a sentence or paragraph), and an AI generated a continuation that attempts to follow the original style, tone, and intent.
  • AI-paraphrased: Texts that were originally written by a human and then reworded by an AI system to express the same meaning using different phrasing, possibly altering sentence structure or word choice while preserving the original message.
By mixing prompts across these categories, we generate a balanced corpus that mitigates over-fitting to any one prompt pattern and better reflects the diversity of real user queries.
Category Student Thesis Paper Abstract
AI-generated You are an IT student who are writing the graduation thesis. In a formal academic tone, write this section of a thesis on       , clearly outlining the structure, then write the passage concisely. The original text:       . In clear, structured prose, draft the section for a thesis titled       , cite some related works you mentioned in the passage and highlighting the contribution. The original text:       . You are a computer scientist who are very familiar with abstract writing for your works based on the title. Craft a concise word_count-word abstract for a paper titled       , summarizing the problem statement, methodology, key findings, and contributions. The original text:       . Compose a word_count-word abstract for the paper       , ensuring it includes motivation, approach, results, and implications for future research. The original text:       .
AI-polished You are an IT student who are refining your work to make it more complete. Refine the following section excerpt for grammar, clarity, and academic style while preserving its original meaning and terminology. The original text:       . You are an IT student who are refining your work to make it more complete. Enhance the academic tone, coherence, and logical flow of this thesis section without altering technical content. The original text:       . You are a computer scientist who are very familiar with abstract writing and refining the written abstract. Improve the coherence, precision, and formal tone of this draft abstract without introducing new content. The original text:       . Improve the clarity and conciseness of this abstract paragraph while maintaining all original findings and terminology. The original text:       .
AI-continued You are an IT student who are writing your graduation thesis. Continue to write the section from this thesis excerpt for approximately word_count words, maintaining formal academic structure and style. The original text:       . You are an IT student who are writing your graduation thesis. Extend the section by adding supporting detailed information for a thesis on       . The original text:       . You are a computer scientist who are very familiar with abstract writing. Add some concise concluding sentences to this partial abstract that highlights implications for future research. The original text:       . Continue the abstract by writing a closing statement that underscores the study’s contributions and potential applications. The original text:       .
AI-paraphrased Paraphrase the following thesis section in a clear academic tone, preserving citations and technical terms exactly. The original text:       . Reword this thesis excerpt to improve readability and maintain its scholarly voice, keeping all references unchanged. The original text:       . Rephrase this abstract in formal academic English, maintaining all original citations and technical accuracy. The original text:       . Paraphrase this abstract paragraph to enhance clarity and flow, ensuring all technical terms and citations remain intact. The original text:       .
Table 6: Samples of diverse prompt templates used to generate FAIDSet
Appendix C Analysis on similarity across models and model families
To assess the stylistic and semantic consistency of AI-generated texts, we conducted a comprehensive analysis using various perspectives: N-grams distribution, text length patterns, and semantic embedding visualization. These analyses allow us to examine similarities both within the same model family and across different model families, leading to a robust understanding of LLM "authorship" characteristics.
C.1 Text Distribution Between AI Families
We first examined the distribution of text length, measured in both word and character counts across outputs from five AI models: Llama-3.3-70B-Instruct-Turbo Dubey et al. (2024), GPT-4o-mini OpenAI (2024), Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash Team et al. (2024). Using 2000 arXiv prompt seeds, each model generated a corresponding output, and their length distributions were plotted.
From Figure 3, distinct patterns emerged between families. Gemini models consistently produced shorter and more compact outputs, while Llama and GPT models demonstrated more variance and a tendency toward longer completions. Despite using different versions (e.g., Gemini 2.0 vs. Gemini 1.5 Flash), the Gemini outputs remained closely grouped in terms of both word and character counts, suggesting a shared output generation strategy and stylistic consistency across the family. In contrast, Llama and GPT distributions showed greater separation and variability.
This reinforces the hypothesis that text length behavior is not only model-dependent but also family-coherent, with Gemini models forming a distinct cluster.
C.2 Text Distribution Between AI Models Within the Same Family
We performed N-grams frequency analysis on responses generated by three models within the Gemini family: Gemini 2.0, Gemini 2.0 Flash, and Gemini 1.5 Flash using 500 texts from the arXiv abstract dataset. Figure 5 highlights overlapping high-frequency tokens and similar patterns in word usage and phrase structure among the three models.
Despite minor differences in architectural speed (e.g., Flash vs. regular) or release chronology, the N-grams distributions show minimal divergence. Frequently used tokens, such as domain-specific terms and transitional phrases, appeared with nearly identical frequencies. This suggests that these models share similar decoding strategies and training biases, likely due to shared pretraining corpora and optimization techniques resulting in highly consistent stylistic patterns.
These intra-family similarities support treating model variants within a family as a unified authoring entity when performing analysis or authorship attribution.
C.3 Embedding Visualization and Semantic Cohesion
To delve into deeper semantic alignment, we visualized embeddings generated by unsupervised SimCSE XLM-RoBERTa-base model of generated texts from two model families: Gemini family (texts generated by Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash), and GPT-4o/4o-mini using PCA to reduce the high-dimensional embeddings into a space.
As shown in Figure 4, Gemini model embeddings form tight, overlapping clusters, indicating a high degree of semantic cohesion among their outputs. This clustering remains consistent across both sample sizes of 2000 texts. In contrast, GPT-4o/4o-mini embeddings occupy a separate region of the space, with greater dispersion and less overlap with Gemini clusters.
This visualization confirms that the Gemini family not only shares stylistic features but also maintains semantic coherence, distinguishing it from models of other families at a conceptual level.
C.4 Conclusion
The consistency observed across N-grams distributions, text length patterns, and semantic embeddings among Gemini models substantiates our decision to treat each AI family as an author. These models demonstrate coherent writing styles, shared lexical preferences, and tightly clustered semantic representations - hallmarks of unified authorship. Conversely, inter-family comparisons show clear separability, emphasizing the distinctiveness of each AI model family’s writing behavior.
(a) Word length distribution across five AI models.
(b) Character length distribution across five AI models.
Figure 3: Text length distributions in words and characters across Llama-3.3, GPT-4o/4o-mini, Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash.
(a) 2D embedding space.
(b) 3D embedding space.
Figure 4: Visualizations showing clustering behavior of Gemini model family (Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash) and GPT-4o/4o-mini using 2D and 3D embeddings with sample size of 2000 texts.
(a) Top 3-grams of Gemini 2.0 (500 samples).
(b) Top 3-grams of Gemini 2.0 Flash (500 samples).
(c) Top 3-grams of Gemini 1.5 Flash (500 samples).
Figure 5: Top 20 most common 3-gram from Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash using 500 sample prompts.
Appendix D Model selection for detector
Model #. Params Known Generators Unseen Generators
Acc $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$ Acc $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
RoBERTa-base 125M 80.09 76.22 0.7328 0.3778 73.45 69.10 0.8901 0.4320
FLAN-T5-base 248M 80.19 75.77 0.7783 0.3947 72.80 68.55 0.9123 0.4467
e5-base-v2 109M 81.53 77.90 0.8023 0.4086 74.21 70.15 0.8804 0.4392
Multilingual-e5-base 278M 91.41 90.82 0.3436 0.1732 85.32 84.50 0.5102 0.2543
XLM-RoBERTa-base 279M 91.90 90.63 0.2345 0.1190 86.75 85.20 0.4125 0.2104
Sup-SimCSE-RoBERTa-base 279M 81.22 78.88 0.7102 0.3619 74.00 71.30 0.8420 0.4251
UnSup-SimCSE-RoBERTa-base 279M 82.19 79.38 0.7156 0.3637 75.10 72.40 0.8305 0.4207
UnSup-SimCSE-XLM-RoBERTa-base 279M 92.12 91.75 0.1904 0.0958 87.45 86.90 0.3507 0.1802
Table 7: Model selection on known vs. unseen generators. The best results in each column are in bold.
To identify the best encoder for our classification task, we evaluated each candidate model on both known generators (FAIDSet testing data) and the new unseen-generator test set introduced in Section 3. Each transformer-based encoder was fine-tuned on FAIDSet training data and then used to predict labels on the two evaluation splits. Table 7 summarizes the accuracy, F1-macro, Mean Squared Error (MSE), and Mean Absolute Error (MAE) for each model under both conditions.
Base model comparison.
We first evaluated three popular monolingual models: RoBERTa-base, Flan-T5-base, and e5-base-v2 using the same training and evaluation splits. RoBERTa-base Liu et al. (2019b) and e5-base-v2 Wang et al. (2022) achieved the most balanced trade-off between classification accuracy and regression error (MSE, MAE), while Flan-T5-base Chung et al. (2022) lagged slightly in F1-macro. These results indicated that a stronger encoder backbone yields more robust performance and motivated exploration of multilingual variants for further gains.
Multilingual variants.
Next, we experimented with XLM-RoBERTa-base Conneau et al. (2019) and Multilingual-e5-base Wang et al. (2024a). Both models leverage cross-lingual pretraining, which in our scenario appears to improve representation of diverse linguistic patterns present in FAIDSet. In particular, XLM-RoBERTa-base delivered a substantial boost in all metrics, suggesting that its multilingual training enhances generalization even on primarily monolingual inputs.
Contrastive learning with SimCSE.
Finally, we integrated contrastive learning via SimCSE Gao et al. (2021) to further refine the sentence embeddings. We tested both supervised (trained on NLI data) and unsupervised (trained on Wikipedia corpus) SimCSE variants applied to RoBERTa-base. The unsupervised version outperformed its supervised counterpart, confirming prior findings that unsupervised SimCSE often yields stronger semantic encoders for downstream tasks. Encouraged by this, we adopted the unsupervised SimCSE approach on XLM-RoBERTa-base Vamvas and Sennrich (2023), which achieved the highest accuracy and lowest error rates across all configurations.
Based on these experiments, we selected the unsupervised SimCSE XLM-RoBERTa-base model for our final system.
Algorithm Known Generators Unseen Generators
Accuracy $\uparrow$ F1-macro $\uparrow$ Accuracy $\uparrow$ F1-macro $\uparrow$
k-Nearest Neighbors (KNN) 90.52 90.21 85.37 84.95
k-Means 88.13 87.48 80.22 79.81
Fuzzy k-Nearest Neighbors 95.18 95.05 93.31 93.25
Fuzzy C-Means 92.67 92.31 90.04 89.53
Table 8: Comparison of clustering algorithms on known vs. unseen generators. The best results are shown in bold.
Appendix E Ablation study
E.1 The need to use a vector database
We first used the trained model to classify the text and observed a significant performance decline from 90% when evaluating on unseen domains or text generated by novel models. This drop reflects the model’s limited ability to generalize beyond the distribution of its original training data. To address this, we propose integrating a vector database that stores dense embeddings of all collected examples, both training and incoming unlabeled data.
By indexing and retrieving semantically similar examples at inference time, the vector database serves as a flexible, scalable memory that bridges gaps between training and test distributions, enhancing the classifier’s resilience to domain shifts and unseen generators. Specifically:
  • Robust Domain Adaptation: New inputs are matched against a broad, continuously growing repository of embeddings, allowing the classifier to leverage analogous instances from related domains without full retraining.
  • Generator-Independent Coverage: As novel text generators emerge, their embeddings populate the database; retrieval naturally adapts to new styles or patterns by finding the closest existing vectors.
E.2 Clustering Algorithm Selection
To improve our detector’s robustness against unseen domains and generators, we evaluated four clustering strategies on the encoded vectors stored in our vector database. Each algorithm was tasked with grouping text samples into human-written, AI-generated, and human-AI collaborative categories, using both known-generator data (held-out from training) and entirely unseen generator data. Performance was measured in terms of accuracy and F1-macro score.
We encoded each text sample using our pretrained sequence-classification model’s penultimate layer, then applied clustering within the vector database to assign soft or hard cluster labels corresponding to the three text-origin classes. From the results in Table 8, we draw some key findings:
  • Traditional algorithms show reasonable performance on held-out known generators but degrade notably on unseen generators.
  • Fuzzy C-Means leverages membership degrees to handle overlapping distributions, improving both accuracy and F1-macro by about 4% over k-Means, with smaller degradation on unseen data.
  • Fuzzy KNN combines local neighbor information with fuzzy membership, achieving the best overall performance: 95.2% accuracy and 95.1% F1-macro on known generators, and 93.3% accuracy and 93.3% F1-macro on unseen generators.
Given its superior ability to adapt to novel domains and generators through weighted neighbor voting and soft cluster assignments, we adopt Fuzzy k-Nearest Neighbors as the clustering component in our full architecture.